Kiles14:

Addison earns a fixed hourly rate working as a sales clerk. If she works on a holiday, she earns a different hourly rate than she earns on a regular day. In one week, she earns $188.50 by working 5 hours on a holiday and 16 hours during regular days. A different week, she earns $254.00 by working 8 hours on a holiday and 20 hours during regular days. How much more is Addison’s holiday hourly rate than her regular hourly rate?

shinedivneonebytime:

whats the important information

Kiles14:

i dont know its for math and i just needed help lol

shinedivneonebytime:

ok so if she earns $188.50

shinedivneonebytime:

and works 5 hours a day we should multiply the 16 hours by the 5 hours to get what?

Kiles14:

lemme see...80 :)

shinedivneonebytime:

good

Kiles14:

but then...

shinedivneonebytime:

theirs 20 hours right

Kiles14:

thats not an answer choice so...we need to do something else, right?

shinedivneonebytime:

yea

shinedivneonebytime:

were not done

shinedivneonebytime:

so if theirs $188.50 and $254.00 add those together

shinedivneonebytime:

ok hold up

Kiles14:

ok...

shinedivneonebytime:

do you know ratios

Kiles14:

i do but i don't remeber lmao

shinedivneonebytime:

ok so if this was a ratio it would be money and hours

Kiles14:

yeah :)

Kiles14:

but then im still not sure

shinedivneonebytime:

look at it this way im going to draw it out

Kiles14:

ok

shinedivneonebytime:



Kiles14:

so whats the next step?

Kiles14:

yk what

Kiles14:

we shoulda divided and then subtracted yk

Kiles14:

But thank you!

shinedivneonebytime:

welcome theirs this person named jjthompson he can help ha will be right here

Kiles14:

are you an adult?

shinedivneonebytime:

no i was just trying to help

Kiles14:

im not, i was just wondering cause your smart!

shinedivneonebytime:

thank you im only in 6th grade

shinedivneonebytime:

lol

Kiles14:

bruh

Kiles14:

im just bad at math ig

Kiles14:

im a freshmen smh🤣

shinedivneonebytime:

wowo

shinedivneonebytime:

wow

shinedivneonebytime:

i think jthompson is here

shinedivneonebytime:

hes gonna help you and me

Kiles14:

bett

shinedivneonebytime:

are you a boy cause most boys say bett

shinedivneonebytime:

smh

Kiles14:

no lmao

Kiles14:

im a girl

shinedivneonebytime:

wow same

Kiles14:

i say bett cause alot of my friends say it lol

jimthompson5910:

Let x = holiday hourly rate (dollars per hour) y = regular hourly rate (dollars per hour) She earns $188.50 by working 5 hours on a holiday and 16 hours during regular days. This means we can set up the equation 5x + 16y = 188.50 Do you agree so far?

shinedivneonebytime:

their jjthompson goes hes real good with math

shinedivneonebytime:

btw that question is for you

Kiles14:

oh yeeeeah mb lol

jimthompson5910:

its ok

Acolgrove:

i love the name in that question

jimthompson5910:

A different week, she earns $254.00 by working 8 hours on a holiday and 20 hours during regular days. So we can also form the equation 8x + 20y = 254 Agreed?

Kiles14:

ohh my gahh

Kiles14:

yeeeeeh

Acolgrove:

Addison is my name lol

Kiles14:

beautiful my gal i love it

Acolgrove:

aww thank u thats sweet

shinedivneonebytime:

kiles you got to be serious

Kiles14:

my name is kileyyyy

shinedivneonebytime:

i know but you got to be serious

Kiles14:

lol its all good thoo

shinedivneonebytime:

i mean idk you name but seriousness

Kiles14:

i answered the question already...it was wrong we dont speak on that

jimthompson5910:

So we have this system of equations \(\large \begin{cases}5x + 16y = 188.50\\8x + 20y = 254\end{cases}\) What we can do is multiply both sides of the first equation by 8 to get 8*(5x+16y) = 8*188.50 40x + 128y = 1508 which will update the first equation As for the second equation, we can multiply both sides by -5 to get 8x + 20y = 254 -5*(8x + 20y) = -5*(254) -40x - 100y = -1270 We have this new system of equations \(\large \begin{cases}40x + 128y = 1508\\-40x - 100y = -1270\end{cases}\) What's the next step from here? Hint: look at how the 40x and -40x combine to 0x or just 0

shinedivneonebytime:

wow now thats being a smartie

shinedivneonebytime:

kiley what are you talking about

Kiles14:

okay he is like really smart but kinda too smart for me, i answered the question but...i got it wrong and um...i do apprecite him alot yk...but im kinda stupid being 14 nd allat but idk my math very well theres somet hings i CAN do but mostly i CANT yk

shinedivneonebytime:

no what jjthompson can you help me instead

Kiles14:

I didn't know he was gonna come until after i answered im sorry

jimthompson5910:

As a quick alternative, you can graph each equation and see where they intersect. Desmos has the feature to be able to click on the intersection points and have the coordinates pop up.

jimthompson5910:

It's ok Kiles14

Kiles14:

you can call me kiley, but i am very sorry :(

jimthompson5910:

no worries kiley

shinedivneonebytime:

its fine im in 6th grade i make mistakes to i hate math and report cards comiing up this fridsay

jimthompson5910:

what was your question @shinedivneonebytime

shinedivneonebytime:

Janelle’s family went grocery shopping and bought 80 eggs. Janelle bought 20 eggs. What percentage of the total number of eggs did Janelle buy?

Kiles14:

okay lmao i be passing with C's nd B's sometimes A's but i try :)

jimthompson5910:

You already asked that one

Kiles14:

you two work it out

shinedivneonebytime:

ik but you didnthelp

jimthompson5910:

You found earlier that 20/80 = 0.25 which converts to 25% You effectively already got the answer

shinedivneonebytime:

Mr. Teague bought cards to celebrate the holidays. He sends 40% of the cards to his friends. He sent 56 cards. How many cards did he buy altogether?

jimthompson5910:

x = number of cards sent total 40% of x = 56 0.40x = 56 Solve for x. You'll need to divide both sides by 0.40

shinedivneonebytime:

so 40 divided by 0.40

jimthompson5910:

56 divided by 0.40

shinedivneonebytime:

or 56 divided o.40 equals 140

jimthompson5910:

yes, so 140 is the answer

shinedivneonebytime:

: Residents of a small city voted on whether to allow a developer to build a shopping center. The number of votes in favor of the shopping center was 4,400. The number of votes against the shopping center was 17,600. What percent of the total votes were in favor of building the shopping center?

jimthompson5910:

4400 voted in favor 17600 voted against how many total voted?

shinedivneonebytime:

it has answer chices

shinedivneonebytime:

A) 20 % B) 25 % C) 40 % D) 44 %

jimthompson5910:

add 4400 to 17600 to get what result?

shinedivneonebytime:

22000

jimthompson5910:

yes, now divide the number in favor (4400) over this total (22000)

jimthompson5910:

4400/22000 = ??

shinedivneonebytime:

0.2

jimthompson5910:

yep and then we move the decimal point to the right 2 spaces to get 20%

shinedivneonebytime:

Arianna made 80% of the cupcakes for a bake sale. She made 240 cupcakes. How many cupcakes were made altogether?

shinedivneonebytime:

thank you so much

jimthompson5910:

x = total number of cupcakes 80% of total = amount made for the bake sale 80% of x = 240 0.80*x = 240 To solve for x, divide both sides by 0.80 So you'll need to use your calculator to determine what 240/0.80 is equal to the / means division